OpenStudy (anonymous):

May you please help me...Kinda suck at this :( –2x8) · 3y9 · 2x4 (1 point) 3x12y9 –12x72y9 –12xy21 –12x12y9

zepdrix (zepdrix):

\[\huge \color{green}{-2x^8}\cdot\color{cornflowerblue}{3y^9}\cdot\color{orangered}{2x^4}\]Let's start by multiplying the green and orange terms first. When you multiply terms with similar bases you ADD the exponents.\[\huge \color{green}{-2}\cdot\color{orangered}{2x^4}\color{green}{x^8} \qquad = \qquad \color{green}{-2}\cdot\color{orangered}{2}x^{4+8}\]

zepdrix (zepdrix):

Which leaves us with this, \[\huge \color{green}{-4x^{12}}\cdot\color{cornflowerblue}{3y^9}\]

zepdrix (zepdrix):

Then just multiply the 3 and the 4 :O Hmm maybe it would have been easier to move the constants out front, and multiply those first. I hope I didn't confuse you :3